OpenStudy (anonymous):

I need help with this problem 7+4x<9

OpenStudy (anonymous):

Please someone help me with this problem I have never used open study just this one question. 7+4x<9

OpenStudy (anonymous):

Solve this inequality for x. 7 + 4x < 9 Write your answer as an inequality including x and < or >. Do not use words in your answer. here is the full question

OpenStudy (anonymous):

Subtract 7 from each side and then divide each side by 4.\[x<\frac{1}{2} \]

OpenStudy (anonymous):

did you see the whole question

OpenStudy (anonymous):

Solve this inequality for x. 7 + 4x < 9 Write your answer as an inequality including x and < or >. Do not use words in your answer.

OpenStudy (anonymous):

\[\frac{-7+7+4x}{4}<\frac{9-7}{4} \]\[x<\frac{1}{2} \]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're Welcome.